OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression - x^2-1 x-1 -------- / ----- 2x 4x^2 **Note the / means to divide

OpenStudy (duck1064):

-5

OpenStudy (mertsj):

\[\frac{x^2-1}{2x}\div \frac{x-1}{4x^2}\] \[\frac{(x-1)(x+1)}{2x}\times \frac{4x^2}{x-1}\]

OpenStudy (mertsj):

Cancel the common factors and multiply